A can eat 3 apples in 5 days while B can eat 6 apples in 4 days.In how many days can the two persons consume 21 apples?
A)10
B)8

Answer by fcabanski(1391)

A eats 3 apples in 5 days, or 3/5 apple per day.


B eats 6 apples in 4 days, or 6/4 apples per day.


Together they eat 3/5 + 6/4 = 12/20 + 30/20 = 42/20 = 21/10 apples per day.


They eat 21 apples in 21 / (21/10) = 21/1 * 10/21 = 10 days.